### v1

# High-Temperature Quantum Anomalous Hall Effect in Buckled Honeycomb Antiferromagnets

Abstract
We propose Néel antiferromagnetic (AF) Mott insulators with a buckled honeycomb structure as potential candidates to host a high-temperature AF Chern insulator (AFCI). Using a generalized Kondo lattice model we show that the staggered potential induced by a perpendicular electric field due to the buckling can drive the AF Mott insulator to an AFCI phase. We address the temperature evolution of the Hall conductance and the chiral edge states. The quantization temperature $T_{q}$, below which the Hall conductance is quantized, depends essentially on the strength of the spin-orbit coupling and the hopping parameter, independent of the specific details of the model. The deviation of the Hall conductance from the quantized value $e^{2}/h$ above $T_{q}$ is found to be accompanied by a spectral broadening of the chiral edge states, reflecting a finite life-time, i.e., a decay. Using parameters typical for heavy transition-metal elements we predict that the AFCI can survive up to room temperature. We suggest Sr3CaOs2O9 as a potential compound to realize a high-$T$ AFCI phase.
Introduction. Due to its far-reaching potential applications in topological quantum computation and low-energy-consumption spintronic devices the Chern insulator (CI) state has attracted considerable attention in the past decade [1, 2]. This has led to the discovery of the CI in different classes of systems including thin films of magnetically doped topological insulators [3, 4, 5], thin films of the intrinsic magnetic topological insulator MnBi2Te4 [6, 7, 8], and moiré materials [9, 10]. Despite this remarkable progress, the observation of the CI is limited to temperatures of only a few Kelvins, arising from the material’s negligible charge gap or low magnetic transition temperature.
While the current realization of the CI is limited to ferromagnets, antiferromagnets are far more common and exhibit generally higher transition temperatures, reaching hundreds of Kelvins [11]. In addition, the AF ordering of strongly correlated electrons is known to be accompanied with a noticeable blue shift of the charge gap [12, 13, 14, 15]. This is to be compared with the ferromagnetic ordering on the top and the bottom surfaces of MnBi2Te4 inducing almost no gap in the Dirac states [16, 17, 18, 19]. Thus, there is compelling reason to search for a high-temperature CI in AF Mott materials.
A non-zero Chern number necessitates the time-reversal symmetry to be truly broken [20, 21]. This is inherent in ferromagnets but not guaranteed in antiferromagnets. The effect of the time-reversal transformation on an AF state can be compensated by a lattice group operation. This composite anti-unitary symmetry needs to be broken for the emergence of an AFCI, which is usually achieved by inducing a staggered potential between the spin-up and the spin-down sublattices [20, 21, 22]. The direction of the magnetization on the higher-energy (or the lower-energy) sublattice determines the clockwise or the counter-clockwise propagation direction of the chiral edge states. The AFCI phase is already predicted in various systems [20, 21, 22, 23, 24, 25]. For those involving heavy transition-metal elements a large charge gap is also reported [24, 25]. However, an explicit study of whether the AFCI can persist up to high temperatures remains a crucial open question.
Here, we investigate the Néel AF transition-metal compounds possessing a buckled honeycomb structure, see Fig. 1(a). Using a generalized Kondo lattice model we confirm that a perpendicular electric field can drive the AF Mott insulating state to an AFCI phase. We address the temperature evolution of the Hall conductance and the chiral edge states. The quantization temperature $T_{q}$, below which the Hall conductance is quantized, shows generic behavior depending only on the spin-orbit coupling and the hopping parameter. The deviation of the Hall conductance from the quantized value $e^{2}/h$ above $T_{q}$ is accompanied by a spectral broadening of the chiral edge modes. For heavy transition-metal elements we estimate that $T_{q}$ can reach room temperature. We propose Sr3CaOs2O9 as a potential candidate to realize a high-$T$ AFCI state.
Model Hamiltonian. The low-energy properties of transition-metal compounds are commonly described by the spin-$\mathcal{S}_{\rm tot}$ Heisenberg model [26]
$H=J\sum_{\langle i,j\rangle}\vec{\mathcal{S}}_{i}\cdot\vec{\mathcal{S}}_{j}+\cdots\ ,$ (1)
where $\langle i,j\rangle$ limits $i$ and $j$ to be nearest-neighbor (NN) and the dots stand for terms beyond the isotropic NN interaction. For the Néel AF order, the NN AF interaction ($J>0$) is usually the dominant term.
For the buckled honeycomb structure, see Fig. 1(a) for a side view, applying a perpendicular electric field induces a potential difference between the spin-up and the spin-down sublattices. This induces a shift of charges between the two kinds of sites. We approximate the multi-orbital Hubbard model by a generalized Kondo lattice model. This is well-justified if the charge fluctuations, i.e., the number of holes or double occupancies, do not exceed one per site. The generalized Kondo lattice model is given by (cf. Fig. 1)
$H=$ $+t\sum_{\langle i,j\rangle}\sum_{\alpha}c^{{\dagger}}_{i\alpha}c^{{\vphantom{\dagger}}}_{j\alpha}-2J_{\rm H}\sum_{i}\vec{s}_{i}\cdot\vec{S}_{i}+U\sum_{i}n_{i,\uparrow}n_{i,\downarrow}$
$+J\sum_{\langle i,j\rangle}(\vec{{S}}_{i}\cdot\vec{{S}}_{j}+\vec{{S}}_{i}\cdot\vec{{s}}_{j}+\vec{{s}}_{i}\cdot\vec{{S}}_{j})$
$+\sum_{i}\sum_{\alpha}\delta_{i}^{{\vphantom{\dagger}}}n^{{\vphantom{\dagger}}}_{i\alpha}+{\rm i}\lambda_{\rm SO}\sum_{[i,j]}\sum_{\alpha\beta}\nu_{ij}^{{\vphantom{\dagger}}}c^{{\dagger}}_{i\alpha}\sigma^{z}_{\alpha\beta}c^{{\vphantom{\dagger}}}_{j\beta}\ ,$ (2)
where $c^{{\dagger}}_{i\alpha}$ is the fermion creation operator at the site $i$ with the $z$-component of the spin $\alpha=\uparrow$ or $\downarrow$. The first term is the NN hopping and the second term is the Hund coupling between the electron spin $\vec{s}_{i}$ and the localized spin $\vec{S}_{i}$ with the spin quantum number $S=\mathcal{S}_{\rm tot}-1/2$. We treat one orbital as representative of the charge motion, while the others represent the localized spin. The third term in Eq. (2) is the Hubbard interaction with $n^{{\vphantom{\dagger}}}_{i\alpha}:=c^{{\dagger}}_{i\alpha}c^{{\vphantom{\dagger}}}_{i\alpha}$, and the fourth term is the Heisenberg interaction between the NN spins. We always count the Heisenberg interaction on each lattice bond only once. The fifth term is the staggered sublattice potential giving the onsite energies $\delta$ and $\delta$ to the two sublattices of the honeycomb structure. Figure 1(b) illustrates these different terms for the special case of $\mathcal{S}_{\rm tot}=3/2$.
Figure 1: (a) Side view of the spin-$\mathcal{S}_{\rm tot}$ Heisenberg model with the nearest-neighbor AF interaction $J$ on the buckled honeycomb structure. In the presence of a perpendicular electric field $\vec{E}=E_{0}\hat{z}$ the system can effectively be described by the generalized Kondo lattice model (2) illustrated in (b) for the special case of $\mathcal{S}_{\rm tot}=3/2$.
The last term in Eq. (2) is included to account for the effect of the spin-orbit coupling. It is exactly the Kane-Mele term [27]. The notation $\left[i,j\right]$ restricts $i$ and $j$ to be next-nearest-neighbor, $\sigma^{z}$ stands for the Pauli matrix, and $\nu_{ij}=2/\sqrt{3}(\hat{d}_{1}\times\hat{d}_{2})_{z}=\pm 1$ where $\hat{d}_{1}$ and $\hat{d}_{2}$ are the unit vectors along the two bonds the electron traverses from site $j$ to site $i$. We add a chemical potential $\mu=U/2$ to the Hamiltonian (2) to satisfy the half-filling condition.
We always focus on large values of $U$ and $J_{\rm H}$ corresponding to the Mott regime and fix the Heisenberg interaction in Eq. (2) to $J=4t^{2}/\Delta_{0}$ with $\Delta_{0}:=U+2SJ_{\rm H}$, called the bare Mott gap. This guarantees the Heisenberg model (1) as the low-energy effective model of the Hamiltonian (2) for $\delta=0$, apart from some weak anisotropic interactions originating from the spin-orbit coupling $\lambda_{\rm SO}$. The Hamiltonian (2) generalizes the purely spin model (1) allowing for the study of the charge fluctuations induced by a perpendicular electric field.
The number of methods which allow for a reliable investigation of the topological properties of the strongly correlated Hamiltonian (2) at finite temperature are rare. We employ the dynamical mean-field theory (DMFT) [28] as an established method for strongly correlated systems and use the exact diagonalization (ED) as the impurity solver [28, 29]. We specifically use the real-space realization of the DMFT [30] providing access to the bulk and the edge properties on equal footing [31]. Technical details can be found in Supplemental Materials [32]. We will compare the results for the different number of bath sites in the ED calculations.
[…]
where $\boldsymbol{G}^{{\vphantom{\dagger}}}_{\alpha}\equiv\boldsymbol{G}^{{\vphantom{\dagger}}}_{\alpha}({\rm i}\omega_{n},\vec{k})$, $\epsilon_{\mu\nu\rho}$ is the totally antisymmetric tensor, and summations are implied over the indices $\mu$, $\nu$, and $\rho$, each of which runs over $\mathrm{i}\omega_{n}$, $k_{x}$, and $k_{y}$ [32].
The momentum-resolved spectral function is accessed using the real-frequency Green’s function,
$A_{\alpha,\vec{d}}~(\omega,\vec{k})=-{\frac{1}{\pi}\rm Im}\!\left[\boldsymbol{G}^{{\vphantom{\dagger}}}_{\alpha}(\omega+{\rm i}\eta,\vec{k})\right]_{\vec{d},\vec{d}}$ (4)
where $\vec{d}$ specifies a lattice site in the unit cell. The broadening factor $\eta=0.01t$ is used in the calculations. The local spectral function $A_{\alpha,\vec{d}}~(\omega)$ is computed by integrating over the momentum with the appropriate prefactor. The spectral sum rule is always satisfied with an error less than $10^{-3}$ in all the results that we present.
Figure 2: Phase diagram of temperature vs the alternating sublattice potential controlled by a perpendicular electric field. The colormap represents the value of the Hall conductance $\sigma_{yx}=-\sigma_{xy}$ given by Eq. (High-Temperature Quantum Anomalous Hall Effect in Buckled Honeycomb Antiferromagnets). The Néel temperature $T_{\rm N}$ and the crossover quantization temperature $T_{q}$ are specified. The Hall conductance takes the quantized value $e^{2}/h$ with an error less than $\%1$ below $T_{q}$, characterizing the antiferromagnetic Chern insulator (AFCI). The gray dotted lines separate the metallic region from the insulating regions. The metallic region shrinks rapidly to the two quantum critical points as $T\to 0$. The results are for the model parameters $S=1/2$, $U=12t$, $J_{\rm H}=0.2U$, $J=4t^{2}/\Delta_{0}=0.2\bar{7}t$, and $\lambda_{\rm SO}=0.2t$. We recall that the localized spin $S=1/2$ in Eq. (2) corresponds to the total spin $\mathcal{S}_{\rm tot}=S+1/2=1$ in Eq. (1). The number of bath sites $n_{b}=5$ is used in the ED impurity solver.
Phase diagram. Figure 2 represents the phase diagram of temperature vs the alternating sublattice potential for the model parameters $S=1/2$, $U=12t$, $J_{\rm H}=0.2U$, $J=4t^{2}/\Delta_{0}=0.2\bar{7}t$, and $\lambda_{\rm SO}=0.2t$. Note that the localized spin $S=1/2$ in Eq. (2) corresponds to the total spin $\mathcal{S}_{\rm tot}=1$ in Eq. (1), see Fig. 1. The number of bath sites $n_{b}=5$ is used in the ED. The colormap displays the value of the Hall conductance $\sigma_{yx}=-\sigma_{xy}$. The Néel temperature $T_{\rm N}$ and the crossover quantization temperature $T_{q}$ are specified. The Hall conductance acquires the quantized value $e^{2}/h$ with an error less than $\%1$ below $T_{q}$, characterizing the AFCI. The gray dotted lines separate the metallic region from the insulating regions. The metallic region shrinks rapidly to the two quantum critical points as $T\to 0$.
Figure 3: Local magnetizations $m=|\langle s_{i}^{z}\rangle|$ and $M=|\langle S_{i}^{z}\rangle|$ (a) and the Hall conductance (b) vs $T$. The Néel temperature $T_{\rm N}$ and the crossover quantization temperature $T_{q}$ are specified. (c) Local spectral function averaged over the two sites in the unit cell for the spin component $\alpha$ at different temperatures. The results correspond to the solution with the magnetization on the higher-energy sublattice pointing in the positive $z$ direction. The results are for $S=1/2$, $U=12t$, $J_{\rm H}=0.2U$, $J=4t^{2}/\Delta_{0}=0.2\bar{7}t$, $\lambda_{\rm SO}=0.2t$, and $\delta=7t$. The data are for the number of bath sites $n_{b}=6$ except for the filled gray squares at selective temperatures in (a) and (b) which are for $n_{b}=7$.
For small values of $\delta$ the system is a paramagnetic Mott insulator at high and an AFI at low $T$, which is the well-known physics of the Heisenberg model [34]. Upon increasing $\delta$, the Néel temperature first increases. This is due to the increase in the effective Heisenberg interaction $J_{\rm eff}=4t^{2}\Delta_{0}/(\Delta_{0}^{2}-4\delta^{2})$ between the itinerant spins. For $\delta\gtrsim 5t$ the charge fluctuations start to contribute to the low-energy physics causing a drop in $T_{\rm N}$. For large values of $\delta$ the lower-energy sublattice is almost doubly occupied and the higher-energy sublattice is almost empty of electrons. The itinerant electrons show a very weak local magnetization due to the van Vleck mechanism [35, 32]. Thus, the itinerant electrons are in the band insulator phase above $T_{\rm N}$.
Figure 4: The brick wall representation of the honeycomb structure with the open boundary condition in $x$ and the periodic boundary condition in $y$ direction. The different sites in the $x$ direction are labeled from 0 to $N_{x}-1$. The dashed box specifies the unit cell.
The intermediate values of $\delta$ are of particular interest because the Hall conductance becomes non-zero and approaches the quantized value $e^{2}/h$ as the temperature is lowered. For $\delta=7t$ we plot the local magnetizations $m=|\langle s_{i}^{z}\rangle|$ and $M=|\langle S_{i}^{z}\rangle|$ in Fig. 3(a) and the Hall conductance vs $T$ in Fig. 3(b). Figure 3(c) represents the local spectral function averaged over the two sites in the unit cell for the spin component $\alpha$ near the Fermi energy $\omega=0$ at different temperatures. The data in Fig. 3 are for $n_{b}=6$ bath sites except for the filled gray squares at selective temperatures in panels (a) and (b) which are for $n_{b}=7$. The results for $n_{b}=5$, not included in the figure, also nicely match the results for $n_{b}=6$ and 7. This corroborates the accuracy of the results.
Figure 5: (a) The momentum-resolved spectral function $A_{\alpha,x}(\omega,k_{y})$ for the topological spin component $\alpha=\uparrow$ at the edge $x=0$ at different temperatures. (b) Local spectral function $A_{\uparrow,x}(\omega)$ for $T=0.1t$ at different $x$. The results are for the same model parameters as in Fig. 3. A cylindrical geometry as illustrated in Fig. 4 with $N_{x}=80$ is used. The data are for $n_{b}=6$ bath sites in the ED impurity solver.
Temperature evolution of chiral edge states. The quantized Hall conductance at intermediate values of $\delta$ in Fig. 2 is expected to support chiral edge states. The direction of magnetization on the higher-energy (or the lower-energy) sublattice determines the clockwise or counter-clockwise propagation direction of chiral edge states. We focus on the solution with the magnetization on the higher-energy sublattice pointing in the positive $z$ direction. The other solution can be accessed using the time-reversal transformation.
To investigate the presence and the temperature evolution of the chiral edge states we consider a cylindrical geometry with open boundary condition in the $x$ direction and edges of type armchair. The honeycomb structure is treated as a brick wall labeling the lattice sites in the $x$ direction from 0 to $N_{x}-1$, as illustrated in Fig. 4. At each $x$, there are two nonequivalent lattice sites in the $y$ direction. We consider $N_{x}=80$.
The momentum-resolved spectral function $A_{\uparrow,x=0}(\omega,k_{y})$, averaged over the two nonequivalent lattice sites in the $y$ direction, is plotted in Fig. 5(a) for the same model parameters as in Fig. 3. The results are for the topological spin component $\alpha=\uparrow$. The spin component $\alpha=\downarrow$ has no contribution to the quantized Hall conductance at low-$T$ in Fig. 3(b) and we find it to be gapped in the bulk and at the edges. The chiral edge state in Fig. 5(a) at $T=0.1t$ quickly disappears as the bulk is approached. This can be seen from the momentum-integrated spectral function $A_{\uparrow,x}(\omega)$ given for different values of $x$ in Fig. 5(b). The temperature evolution of the momentum-resolved spectral function in Fig. 5(a) indicates a spectral broadening as the temperature is increased above $T\sim 0.2t$. This reflects a finite life-time, i.e., a decay, of the chiral edge state and is in accord with the deviation of the Hall conductance from the quantized value $e^{2}/h$ above $T_{q}$ in Fig. 3(b).
Figure 6: Quantization temperature of the Hall conductance vs the spin-orbit coupling for different sets of model parameters. The Heisenberg interaction is always fixed to $J=4t^{2}/\Delta_{0}$. In each case, the optimal value of $\delta$ is chosen such that the system is in the middle of the AFCI phase. The results are for the number of bath sites $n_{b}=5$. The dashed line represents a linear fit to the data for $\lambda_{\rm SO}<0.15t$.
Quantization temperature. Up to now, our study has been limited only to particular model parameters. Nevertheless, the AFCI always appears when the sublattice potential difference $2\delta$ is of the same order as the bare Mott gap $\Delta_{0}=U+2SJ_{\rm H}$, independent of the details of the model. The quantization temperature $T_{q}$ also indicates generic behavior. To demonstrate this, we have plotted $T_{q}$ vs the spin-orbit coupling in Fig. 6 for different sets of model parameters. In each case, the optimal value of $\delta$ is chosen such that the system is in the middle of the AFCI phase, see Fig. 2 for $\delta=7.2t$ used in the last set of parameters. The Heisenberg interaction is always fixed to $J=4t^{2}/\Delta_{0}$.
Interestingly, despite the very different model parameters used in Fig. 6 one can hardly see any change in $T_{q}$. The quantization temperature is solely determined by the spin-orbit coupling and the hopping parameter. $T_{q}$ increases linearly with $\lambda_{\rm SO}$ and saturates for large values of $\lambda_{\rm SO}$. Although the spin-orbit coupling is highly compound-specific, one usually expect $\lambda_{\rm SO}<0.15t$. In this region, the data can nicely be described by the linear fit $T_{q}=0.92\lambda_{\rm SO}-0.003t$ shown in Fig. 6 as a black dashed line. This relation suggests a quantization temperature of the order of room temperature for $\lambda_{\rm SO}\sim 30$ meV and $t\sim 500$ meV, which are typical values for heavy transition-metal elements.
Conclusion. We propose the strongly correlated AF materials with a buckled honeycomb structure as potential candidates to realize a hight-$T$ CI phase. The system can be driven from the AF Mott insulating state to the AFCI phase by applying a perpendicular electric field, which allows to induce and fine-tune a sublattice potential difference. The AFCI emerges when the sublattice potential difference reaches the same size as the Mott gap. The experimentally accessible electric field of $0.5$ V/Å [36, 37, 38] can induce a sublattice potential difference of a few eV in a buckling height of a few angstroms. Since the Mott gap and the buckling height can also be effectively controlled by pressure, we believe the AFCI state is realistically within reach.
Compounds Ba2NiTeO6 and Sr3CaOs2O9 are known buckled honeycomb antiferromagnets with the buckling heights of $5.3$ Å [39, 40] and $3.3$ Å [41], respectively. The latter compound shows a Néel AF order with the high Néel temperature $T_{\rm N}\sim 385$ K and is a promising candidate to host a high-$T$ AFCI. Growing $[111]$-oriented bilayers of perovskite heavy transition-metal oxides is another route to a buckled honeycomb antiferromagnet [42] and a high-$T$ AFCI state. Our findings open up a new path to the realization of the quantum anomalous Hall effect at high temperatures.